# Independence Assumption RLS Regularized Least Squares

### v1

@@ 1 Introduction
This observation can be used to measure how far the language model can be from the ground truth. The sequential model implemented by a language model allows us to construct a joint distribution over multiple responses, which is done through iterative prompting of an LLM based on its previous responses and the application of the chain rule of probability: first we ask the model to provide a response given a query, then to provide another response given the query and the first response, then a third one given the query and the first two responses, an so on. This is in contrast to some of the earlier works that approached decoupling epistemic and aleatoric uncertainty for classification problems by training the model with pairs (or tuples) of labels (Wen et al., 2022).
@@
Interestingly, as we will see in Section˜3, we can force an LLM to provide a desired (possibly incorrect) response by adding this response repeatedly to the prompt. This phenomenon is then further investigated from the viewpoint of a transformer LLM architecture in Section˜3.1.
@@
(i) Based on the above iterative prompting procedure, we derive an information-theoretic metric of epistemic uncertainty in LLMs (Section˜4), which quantifies the gap between the LLM-derived distribution over responses and the ground truth. This gap is insensitive to aleatoric uncertainty, and therefore we can quantify epistemic uncertainty even in cases where there are multiple valid responses.
@@
As usual, $\mathbb{N}$ and ${\mathbb{R}}$ denote the sets of natural and real numbers, respectively. For any measurable set $\mathcal{Z}$, we denote the the set of distributions supported on $\mathcal{Z}$ by ${\mathcal{M}}_{1}(\mathcal{Z})$. For any positive integer $k$, we denote $[k]=\{1,\ldots,k\}$.
@@ 2 Preliminaries
Let $\mu,\mu^{\prime}$ be distributions supported on set $\mathcal{Z}=\mathcal{Z}_{1}\times\cdots\times\mathcal{Z}_{n}$ where $(\mathcal{Z}_{i})_{i}$ is a collection of countable sets. The entropy of a distribution $\mu$ is defined as $H(\mu)=\sum_{z\in\mathcal{Z}}\mu(z)\ln(1/\mu(z))$. [^4] Following the usual convention, we define $0\ln 0=0$ and $a\ln(a/0)=\infty$ for any $a>0$. If $\mu,\mu^{\prime}$ are such that $\mu^{\prime}(z)=0$ only if $\mu(z)=0$, we have a Kullback-Leibler (KL) divergence between them defined as $D_{\mathrm{KL}}(\mu,\mu^{\prime})=\sum_{z\in\mathcal{Z}}\mu(z)\ln(\mu(z)/\mu^{\prime}(z))$. For any $z\in\mathcal{Z}$, we denote $z^{\backslash i}=(z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n})$, and the marginal of the $i$th coordinate of $\mu$ is given by $\mu_{i}(z)=\sum_{z^{\backslash i}\in\mathcal{Z}^{n-1}}\mu(z)$. The product distribution of the marginals of $\mu$ is given by $\mu^{\otimes}(z)=\prod_{i=1}^{n}\mu_{i}(z)$, and the mutual information of $\mu$ is defined as $I(\mu)=D_{\mathrm{KL}}(\mu,\mu^{\otimes})$.
@@ 3 Probability amplification by iteratively prompting
In this section we demonstrate that, as mentioned in the introduction, repeating possible responses several times in a prompt can have pronounced effects on the output of a language model. Consider $x=$“What is the capital of the UK?” and $Y_{1}=\cdots=Y_{t}=$“Another answer to question Q is Paris.” Here we can repeat the sentence “Another answer to question Q is Paris.” an arbitrary number of times. Although the number of repetitions changes the behavior of the LLM, the correct response maintains a significant probability: as Figure˜2 shows, the conditional normalized probability [^5] To obtain conditional normalized probabilities, we consider the probabilities of the two responses, and normalize them so that they add to 1. of the correct response, “London”, reduces from approximately 1 to about 96% as we increase the number of repetitions of the incorrect response to 100. Figure˜2 shows 3 more examples where, with initially low epistemic uncertainty in the response to the query (the aleatoric uncertainty is also low as we consider single-response queries), the correct response maintains a significant or non-negligible probability even in the presence of repetitions of incorrect information, while the probability of predicting the latter is increased.
Q: What is the capital of the UK? A: London (1.0$\approx$ 1.0≈ 1.0) and Paris ($1.29\times 10^{-10}$).
Q: Who was the first US president? A: George Washington (0.9990.9990.999) and Abraham Lincoln ($3.1\times 10^{-06}$).
Q: Who is the author of The Grapes of Wrath? A: John Steinbeck (1.0$\approx$ 1.0≈ 1.0) and Ernest Hemingway ($1.34\times 10^{-10}$).
Q: What is the largest country in the world? A: Russia (0.9990.9990.999) and United Kingdom ($9.02\times 10^{-06}$).
@@
Q: What is the national instrument of Ireland? A: The harp (0.9360.9360.936) and Uilleann pipes (0.0630.0630.063).
Q: Which actor became M in the Bond film Skyfall? A: Ralph Fiennes (0.6510.6510.651) and Judi Dench (0.3480.3480.348).
Q: Which can last longer with out water a camel or a rat? A: A rat (0.5380.5380.538) and A camel (0.4610.4610.461).
Q: If Monday’s child is fair of face what is Saturday’s child? A: Work hard for a living (0.0930.0930.093) and Full of grace (0.9060.9060.906).
@@
Next, we consider a queries for which the model is more uncertain. For the prompt “What is the national instrument of Ireland?”, we observe that responses “The harp” and “Uilleann pipes” both have significant probabilities (the first answer is the correct one). This time, by incorporating the incorrect response in the prompt multiple times, the probability of the correct answer quickly collapses to near zero, as shown in Figure˜2, together with three more examples with significant epistemic uncertainty.
Finally, we consider multi-label queries for which the LLM confidently knows a correct answer. This time, by incorporating a potential response in the prompt, the probabilities of other correct answers stay relatively large. Figure˜3 shows four such examples.
Q: Name a city in the UK A: London (0.9580.9580.958) and Manchester (0.0410.0410.041).
Q: Name a yellow fruit A: Banana (0.7150.7150.715) and Lemon (0.2840.2840.284).
Q: Name an alcoholic drink, A: Wine (0.6850.6850.685) and Beer (0.3140.3140.314).
Q: Name a ball game that is played by more than 5 players A: Volleyball (0.5420.5420.542) and Soccer (0.4570.4570.457).
@@
We consider an idealized attention mechanism as follows. Let ${\mathbf{Z}}\in{\mathbb{R}}^{n\times d^{\prime}}$ be an input matrix comprised of $n$ semantic feature vectors each of dimension $d^{\prime}$. Each row is meant to represent a complete statement (such as “What is the capital of the UK?” or “One answer to the question is Paris.”, etc.) rather than a single token. Let $X^{\top}\in{\mathbb{R}}^{1\times d^{\prime}}$ be the first row of ${\mathbf{Z}}$, which represents the query of interest, such as “What is the capital of the UK?”. Let $E^{\top}\in{\mathbb{R}}^{1\times d^{\prime}}$ be a special vector indicating the end of the input. The matrix ${\mathbf{Z}}\setminus X$, denoting the ${\mathbf{Z}}$ matrix without its first row, represents the in-context information.
@@
Let ${\mathbf{W^{Q}}},{\mathbf{W^{K}}},{\mathbf{W^{V}}}\in{\mathbb{R}}^{d^{\prime}\times d}$ be the query, key, and value matrices. A self-attention head with query $X$ and context ${\mathbf{Z}}\setminus X$ is defined as
@@
If $X$ has appeared many times in the training data, then parameters ${\mathbf{W^{Q}}}$ and ${\mathbf{W^{K}}}$ could be learned such that $E^{\top}{\mathbf{W^{Q}}}({\mathbf{W^{K}}})^{\top}X$ is large, that is, $X$ is within the space spanned by the large principal components of the key-query matrix product. Then, no matter what in-context information appears in ${\mathbf{Z}}$, the probability assigned to $X$ will dominate the sotmax, and we will have
@@ 4 Metric of epistemic uncertainty and its estimation
In this section we apply iterative prompting to estimate the epistemic uncertainty of a language model about responding to some query. The idea is to utilize the different behavior patterns observed in Section˜3, which can be used to differentiate between two modes of high uncertainty: when the aleatoric uncertainty is high vs. when only the epistemic uncertainty is high. We then apply our new uncertainty metric to design a score-based hallucination detection algorithm.
Recall the family of prompts $\mathcal{F}$ defined in Section˜2. We make the following assumption about the ground truth, which states that multiple responses to the same question drawn according to the ground truth are independent from each other:
@@
Here $D_{\mathrm{KL}}$ measures how well $\widetilde{Q}$ approximates $\widetilde{P}$ for a given query $x$. Namely, this metric determines if $\widetilde{Q}$ assigns a large probability to an event which has a small probability under $\widetilde{P}$. In case of LLMs, this means the LLM generates a sequence that is unlikely in the typical usage of the language. In Figure˜4 we have a situation where $\widetilde{P}$ is a pseudo joint distribution derived from the ground-truth, and $\widetilde{Q}$ suffers from a high hallucination rate. Given an input $x$, we want to estimate the above hallucination metric, but we only have access to $\widetilde{Q}$, and so computing it explicitly is impossible. However, next we show that under ˜4.1 we can lower bound $D_{\mathrm{KL}}(\widetilde{Q},\widetilde{P})$ by a quantity which only depends on $\widetilde{Q}$ (the proof is given in Appendix˜B).
@@
For all pseudo joint distributions $\widetilde{P}$ satisfying ˜4.1, we have that
@@
Also, note that we have $I(\widetilde{Q})=D_{\mathrm{KL}}(\widetilde{Q},\widetilde{Q}^{\otimes})$ , $\widetilde{Q}^{\otimes}=\prod_{i}\sum_{y^{\backslash i}}\widetilde{Q}(y_{1},\ldots,y_{i-1},Y_{i},y_{i+1},\ldots,y_{n})$. In general $\sum_{y^{\backslash i}}\widetilde{Q}(y_{1},\ldots,y_{i-1},Y_{i},y_{i+1},\ldots,y_{n})\not=\widetilde{Q}(Y_{i})$, because the independence assumption ˜4.1 does not necessarily (and, in practice, almost never) holds for $Q$.
@@ 4.1 A computable lower bound on epistemic uncertainty
Theorem˜4.5 gives a lower bound on the epistemic uncertainty by the mutual information. However, to compute the mutual information term, in practice we need to evaluate $\widetilde{Q}$ on its entire support, which is potentially infinite. Practically speaking, it is impossible to observe probabilities of all strings under the language model and so we must rely on a finite sample. Therefore, we replace $\widetilde{Q}$ with an empirical distribution with a finite support; in the following we show that the error induced by such an approximation is controlled. To estimate the MI we employ the method given in Algorithm˜1; for generality it is presented for an arbitrary (pseudo) joint distribution $\mu$, but we keep in mind that our case of interest is $\mu=\widetilde{Q}$.
Algorithm 1 MI estimator
[…]
@@
3: Construct a set of indices of unique elements $S=\big{\{}i\in[k]~:~X_{i}\neq X_{j}\quad\forall j<i\big{\}}$
4: Construct empirical distributions: for all $i\in S$,
$\widehat{\mu}(X_{i})$ $=\frac{\mu(X_{i})}{Z}~,$
$\widehat{\mu}^{\otimes}(x^{\prime})$ $=\prod_{i=1}^{n}\frac{\sum_{x^{\backslash i}}\mu(x_{1},\ldots,x_{i-1},x_{i}^{\prime},x_{i+1},\ldots,x_{n})}{Z^{\otimes}}\qquad\text{(for all $x^{\prime}\in{\mathcal{X}}^{n}$)}$
where $Z=\sum_{j\in S}\mu(X_{j})~,\quad Z^{\otimes}=\sum_{j\in S}\sum_{x^{\backslash i}}\mu(x_{1},\ldots,x_{i-1},X_{j,i},x_{i+1},\ldots,x_{n})$
@@
$\widehat{I}_{k}(\gamma)=\sum_{i\in S}\widehat{\mu}(X_{i})\,\ln\left(\frac{\widehat{\mu}(X_{i})+\gamma/Z}{\widehat{\mu}^{\otimes}(X_{i})+\gamma/Z^{\otimes}}\right)$
Note that the estimator is constructed using only each unique element in the sample (the indices of these representative elements are collected in $S$), that is, we do not account for duplicate samples. [^6] When the $X_{i}$ are responses from language models (i.e., when we consider the responses $X_{i}=(Y_{1},\ldots,Y_{n})$), we usually consider the equality of $X_{i}=X_{j}$ semantically, as defined by some function comparator function aiming to compare too responses semantically (our choices for the experiments are described in Section 6). Also note that most terms in the summations defining the product distribution $\widehat{\mu}^{\otimes}$ are zero (except the ones which correspond to the observed data). Furthermore, adding $\gamma/Z$ and $\gamma/Z^{\otimes}$ in the estimator $\widehat{I}_{k}(\gamma)$ is intended to account for the total probability of missing observations, not included while constructing $\widehat{\mu}$ and $\widehat{\mu}^{\otimes}$, making sure the estimate is bounded. Similar ideas are well-know in probability and information theory, such as in universal coding (cesa2006prediction), Laplace smoothing (Polyanskiy and Wu, 2024) and Good-Turing smoothing (Gale and Sampson, 1995, McAllester and Schapire, 2000).
The bias introduced by $\gamma$ in the last equation allows us to rigorously bound the error in estimating $I(\mu)$ via $\widehat{I}_{k}(\gamma)$, which is explored next. In particular, in Theorem˜4.6 we prove a high-probability lower bound on $I(\mu)$ in terms of $\widehat{I}_{k}$. The core of controlling the estimation error is in accounting for the missing mass, or in other words, how much of $\mu$ we miss out by only observing a finite sample. In Appendix˜C, we present a more complete discussion and the proof of the bound on the estimation error for mutual information. Here we adapt this result to our particular case.
@@
Using this quantity, we are ready to present a non-asymptotic bound on the estimation error, which depends on the estimator $\widehat{I}_{k}(\gamma)$, the expected missing mass, and the sample size:
@@
Suppose that $\widehat{I}_{k}(\gamma)$ is given by Algorithm˜1, and assume that ${\mathcal{X}}$ is finite. For $\gamma=1/(k\,|{\mathcal{X}}^{n}|)$, with probability at least $1-\delta$,
$I(\mu)\geq(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma)-\left(\frac{1}{k}+(1+n\,\ln\big{(}1+k\,|{\mathcal{X}}|)\big{)}\,\varepsilon_{k}\right)$
@@
Furthermore, given $\delta_{\mathrm{supp}}\in[0,1)$, let $\tilde{\mathcal{X}}\subseteq{\mathcal{X}}^{n}$ such that $\mu(\tilde{\mathcal{X}})\geq 1-\delta_{\mathrm{supp}}$. Then, for $\gamma=1/(k\,|\tilde{\mathcal{X}}|)$, with probability at least $1-\delta$,
$I(\mu)\geq(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma)-\left(\frac{1}{k}+(1+\ln\big{(}1+k\,|\tilde{\mathcal{X}}|)\big{)}\left(\delta_{\mathrm{supp}}+\varepsilon_{k}\right)\right)~.$
The theorem is a corollary of Theorem˜C.4 shown in Appendix˜C. Note that in Theorem˜4.6 we consider two bounds. The first one is pessimistic in the sense that it does not expect that the samples carry much information about the support, and it is most suitable in situations where we expect $\mu$ to be spread out (uniformly) across its entire support. The price of not having samples covering the whole support in this case is a factor $n\ln|{\mathcal{X}}|$ appearing in the bound. For example, in case of a language model with 10,00010,00010 , 000 tokens, considering all possible strings of length $T$ tokens yields $n\ln|{\mathcal{X}}|=n\,T\ln(10000)$, and so
$I(\mu)\geq(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma)-\left(\frac{1}{k}+(1+n\,T\ln\big{(}1+k\,\ln(10000))\big{)}\,\varepsilon_{k}\right)~.$
Arguably, in practice, such situations are rare, as in natural languages we will not encounter all possible strings. To this end, we consider an optimistic scenario where the effective support of $\mu$, denoted by $\tilde{\mathcal{X}}$, is small with high probability. In this case, we can replace the size of the support for strings of length $n$, $|{\mathcal{X}}|^{n}$, in the first bound with the effective support size $|\tilde{\mathcal{X}}|$, and we only pay essentially a factor $\ln(1+k|\tilde{\mathcal{X}}|)$ instead of $n\ln(1+k|{\mathcal{X}}|)$. In case the effective sample size is only polynomial in $n$, this leads to an exponential reduction in $n$ for the second term in the bounds. In fact, in Section˜C.4 we demonstrate some empirical evidence that on two question-answering benchmarks, $|\tilde{\mathcal{X}}|$ rarely exceeds 100$\approx$ 100≈ 100 with $\mu(\tilde{\mathcal{X}})\geq 0.95$, while sampling responses from an LLM given a query.

### v2

@@ 1 Introduction
This observation can be used to measure how far the language model can be from the ground truth. The sequential model implemented by a language model allows us to construct a joint distribution over multiple responses, which is done through iterative prompting of an LLM based on its previous responses and the application of the chain rule of probability: first we ask the model to provide a response given a query, then to provide another response given the query and the first response, then a third one given the query and the first two responses, an so on. This is in contrast to some of the earlier works that approached decoupling epistemic and aleatoric uncertainty for classification problems by training the model with pairs (or tuples) of labels (Wen et al., 2022, Johnson et al., 2024).
@@
Interestingly, as we will see in Section 3, we can force an LLM to provide a desired (possibly incorrect) response by adding this response repeatedly to the prompt. This phenomenon is then further investigated from the viewpoint of a transformer LLM architecture in Section 3.1.
@@
(i) Based on the above iterative prompting procedure, we derive an information-theoretic metric of epistemic uncertainty in LLMs (Section 4), which quantifies the gap between the LLM-derived distribution over responses and the ground truth. This gap is insensitive to aleatoric uncertainty, and therefore we can quantify epistemic uncertainty even in cases where there are multiple valid responses.
@@
As usual, $\mathbb{N}$ and denote the sets of natural and real numbers, respectively. For any measurable set $\mathcal{Z}$, we denote the the set of distributions supported on $\mathcal{Z}$ by ${\mathcal{M}}_{1}(\mathcal{Z})$. For any positive integer $k$, we denote $[k]=\{1,\ldots,k\}$.
@@ 2 Preliminaries
Let $\mu,\mu^{\prime}$ be distributions supported on set $\mathcal{Z}=\mathcal{Z}_{1}\times\cdots\times\mathcal{Z}_{n}$ where $(\mathcal{Z}_{i})_{i}$ is a collection of countable sets. The entropy of a distribution $\mu$ is defined as $H(\mu)=\sum_{z\in\mathcal{Z}}\mu(z)\ln(1/\mu(z))$. [^4] Following the usual convention, we define $00\ln 0=$ 0 and $a\ln(a/0)=\infty$ for any $a>0$. If $\mu,\mu^{\prime}$ are such that $\mu^{\prime}(z)=0$ only if $\mu(z)=0$, we have a Kullback-Leibler (KL) divergence between them defined as $D_{\mathrm{KL}}(\mu,\mu^{\prime})=\sum_{z\in\mathcal{Z}}\mu(z)\ln(\mu(z)/\mu^{\prime}(z))$. For any $z\in\mathcal{Z}$, we denote $z^{\backslash i}=(z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n})$, and the marginal of the $i$th coordinate of $\mu$ is given by $\mu_{i}(z)=\sum_{z^{\backslash i}\in\mathcal{Z}^{n-1}}\mu(z)$. The product distribution of the marginals of $\mu$ is given by $\mu^{\otimes}(z)=\prod_{i=1}^{n}\mu_{i}(z)$, and the mutual information of $\mu$ is defined as $I(\mu)=D_{\mathrm{KL}}(\mu,\mu^{\otimes})$.
@@ 3 Probability amplification by iteratively prompting
In this section we demonstrate that, as mentioned in the introduction, repeating possible responses several times in a prompt can have pronounced effects on the output of a language model. Consider $x=$“What is the capital of the UK?” and $Y_{1}=\cdots=Y_{t}=$“Another answer to question Q is Paris.” Here we can repeat the sentence “Another answer to question Q is Paris.” an arbitrary number of times. Although the number of repetitions changes the behavior of the LLM, the correct response maintains a significant probability: as Figure 2 shows, the conditional normalized probability [^5] To obtain conditional normalized probabilities, we consider the probabilities of the two responses, and normalize them so that they add to 1. of the correct response, “London”, reduces from approximately 1 to about 96% as we increase the number of repetitions of the incorrect response to 100. Figure 2 shows 3 more examples where, with initially low epistemic uncertainty in the response to the query (the aleatoric uncertainty is also low as we consider single-response queries), the correct response maintains a significant or non-negligible probability even in the presence of repetitions of incorrect information, while the probability of predicting the latter is increased.
Q: What is the capital of the UK? A: London ($\approx 1.0$) and Paris ($1.29\times 10^{-10}$).
Q: Who was the first US president? A: George Washington ($0.999$) and Abraham Lincoln ($3.1\times 10^{-06}$).
Q: Who is the author of The Grapes of Wrath? A: John Steinbeck ($\approx 1.0$) and Ernest Hemingway ($1.34\times 10^{-10}$).
Q: What is the largest country in the world? A: Russia ($0.999$) and United Kingdom ($9.02\times 10^{-06}$).
@@
Q: What is the national instrument of Ireland? A: The harp ($0.936$) and Uilleann pipes ($0.063$).
Q: Which actor became M in the Bond film Skyfall? A: Ralph Fiennes ($0.651$) and Judi Dench ($0.348$).
Q: Which can last longer with out water a camel or a rat? A: A rat ($0.538$) and A camel ($0.461$).
Q: If Monday’s child is fair of face what is Saturday’s child? A: Work hard for a living ($0.093$) and Full of grace ($0.906$).
@@
Next, we consider a queries for which the model is more uncertain. For the prompt “What is the national instrument of Ireland?”, we observe that responses “The harp” and “Uilleann pipes” both have significant probabilities (the first answer is the correct one). This time, by incorporating the incorrect response in the prompt multiple times, the probability of the correct answer quickly collapses to near zero, as shown in Figure 2, together with three more examples with significant epistemic uncertainty.
Finally, we consider multi-label queries for which the LLM confidently knows a correct answer. This time, by incorporating a potential response in the prompt, the probabilities of other correct answers stay relatively large. Figure 3 shows four such examples.
Q: Name a city in the UK A: London ($0.958$) and Manchester ($0.041$).
Q: Name a yellow fruit A: Banana ($0.715$) and Lemon ($0.284$).
Q: Name an alcoholic drink, A: Wine ($0.685$) and Beer ($0.314$).
Q: Name a ball game that is played by more than 5 players A: Volleyball ($0.542$) and Soccer ($0.457$).
@@
We consider an idealized attention mechanism as follows. Let ${\mathbf{Z}}\in^{n\times d^{\prime}}$ be an input matrix comprised of $n$ semantic feature vectors each of dimension $d^{\prime}$. Each row is meant to represent a complete statement (such as “What is the capital of the UK?” or “One answer to the question is Paris.”, etc.) rather than a single token. Let $X^{\top}\in^{1\times d^{\prime}}$ be the first row of ${\mathbf{Z}}$, which represents the query of interest, such as “What is the capital of the UK?”. Let $E^{\top}\in^{1\times d^{\prime}}$ be a special vector indicating the end of the input. The matrix ${\mathbf{Z}}\setminus X$, denoting the ${\mathbf{Z}}$ matrix without its first row, represents the in-context information.
@@
Let ${\mathbf{W^{Q}}},{\mathbf{W^{K}}},{\mathbf{W^{V}}}\in^{d^{\prime}\times d}$ be the query, key, and value matrices. A self-attention head with query $X$ and context ${\mathbf{Z}}\setminus X$ is defined as
@@
If $X$ has appeared many times in the training data, then parameters ${\mathbf{W^{Q}}}$ and ${\mathbf{W^{K}}}$ could be learned such that $E^{\top}{\mathbf{W^{Q}}}({\mathbf{W^{K}}})^{\top}X$ is large, that is, $X$ is within the space spanned by the large principal components of the key-query matrix product. Then, no matter what in-context information appears in ${\mathbf{Z}}$, the probability assigned to $X$ will dominate the softmax, and we will have
@@ 4 Metric of epistemic uncertainty and its estimation
In this section we apply iterative prompting to estimate the epistemic uncertainty of a language model about responding to some query. The idea is to utilize the different behavior patterns observed in Section 3, which can be used to differentiate between two modes of high uncertainty: when the aleatoric uncertainty is high vs. when only the epistemic uncertainty is high. We then apply our new uncertainty metric to design a score-based hallucination detection algorithm.
We will first present the uncertainty metric and its estimate for a distribution defined on the direct outputs of an LLM, and then in Section 4.2, we discuss the changes needed to take semantic equivalences of language into account (Kuhn et al., 2023, Farquhar et al., 2024).
Recall the family of prompts $\mathcal{F}$ defined in Section 2. We make the following assumption about the ground truth, which states that multiple responses to the same question drawn according to the ground truth are independent from each other:
@@
Here $D_{\mathrm{KL}}$ measures how well $\widetilde{Q}$ approximates $\widetilde{P}$ for a given query $x$. Namely, this metric determines if $\widetilde{Q}$ assigns a large probability to an event which has a small probability under $\widetilde{P}$. In case of LLMs, this means the LLM generates a sequence that is unlikely in the typical usage of the language. In Figure 4 we have a situation where $\widetilde{P}$ is a pseudo joint distribution derived from the ground-truth, and $\widetilde{Q}$ suffers from a high hallucination rate. Given an input $x$, we want to estimate the above hallucination metric, but we only have access to $\widetilde{Q}$, and so computing it explicitly is impossible. However, next we show that under 4.1 we can lower bound $D_{\mathrm{KL}}(\widetilde{Q},\widetilde{P})$ by a quantity which only depends on $\widetilde{Q}$ (the proof is given in Appendix D).
@@
For all pseudo joint distributions $\widetilde{P}$ satisfying 4.1, we have that
@@
Also, note that we have $I(\widetilde{Q})=D_{\mathrm{KL}}(\widetilde{Q},\widetilde{Q}^{\otimes})$ , $\widetilde{Q}^{\otimes}=\prod_{i}\sum_{y^{\backslash i}}\widetilde{Q}(y_{1},\ldots,y_{i-1},Y_{i},y_{i+1},\ldots,y_{n})$. In general $\sum_{y^{\backslash i}}\widetilde{Q}(y_{1},\ldots,y_{i-1},Y_{i},y_{i+1},\ldots,y_{n})\not=\widetilde{Q}(Y_{i})$, because the independence assumption 4.1 does not necessarily (and, in practice, almost never) holds for $Q$.
@@ 4.1 A computable lower bound on epistemic uncertainty
Theorem 4.5 gives a lower bound on the epistemic uncertainty by the mutual information. However, to compute the mutual information term, in practice we need to evaluate $\widetilde{Q}$ on its entire support, which is potentially infinite. Practically speaking, it is impossible to observe probabilities of all strings under the language model and so we must rely on a finite sample. Therefore, we replace $\widetilde{Q}$ with an empirical distribution with a finite support; in the following we show that the error induced by such an approximation is controlled. To estimate the MI we employ the method given in Algorithm 1; for generality it is presented for an arbitrary (pseudo) joint distribution $\mu$, but we keep in mind that our case of interest is $\mu=\widetilde{Q}$.
Algorithm 1 MI estimator.
1:  Input: $\mu\in{\mathcal{M}}_{1}({\mathcal{X}}^{n})$ any (pseudo-) joint distribution over ${\mathcal{X}}^{n}$ $k\in\mathbb{N}$ sample size $\gamma_{1},\gamma_{2}\geq 0$ stabilization parameters (typically selected as $1/k$)
@@
3:  Construct a set of indices of unique elements $U=\big{\{}i\in[k]~{}:~{}X_{i}\neq X_{j}\quad\forall j<i\big{\}}$
4:  Construct empirical distributions: for all $i\in U$,
$\widehat{\mu}(X_{i})$ $=\frac{\mu(X_{i})}{Z}~{},\quad\text{where}\quad Z=\sum_{j\in U}\mu(X_{j})$
$\widehat{\mu}^{\otimes}(X_{i})$ $=\prod_{j=1}^{n}\sum_{t\in U:X_{t,j}=X_{i,j}}\hat{\mu}(X_{t,1},\ldots,X_{i,j},\ldots,X_{t,n})$
@@
$\widehat{I}_{k}(\gamma_{1},\gamma_{2})=\sum_{i\in U}\widehat{\mu}(X_{i})\,\ln\left(\frac{\widehat{\mu}(X_{i})+\gamma_{1}}{\widehat{\mu}^{\otimes}(X_{i})+\gamma_{2}}\right)$
Note that most terms in the summations defining the product distribution $\widehat{\mu}^{\otimes}$ are zero (except the ones which correspond to the observed data). Adding $\gamma_{1}$ and $\gamma_{2}$ in the estimator $\widehat{I}_{k}(\gamma_{1},\gamma_{2})$ is intended to account for the total probability of missing observations, not included while constructing $\widehat{\mu}$ and $\widehat{\mu}^{\otimes}$, making sure the estimate is bounded. Similar ideas are well-know in probability and information theory, such as in universal coding (Cesa-Bianchi and Lugosi, 2006), Laplace smoothing (Polyanskiy and Wu, 2024) and Good-Turing smoothing (Gale and Sampson, 1995, McAllester and Schapire, 2000). In Section 4.2, we show an extension of the algorithm that takes semantic equivalences into account, and in the experiments section, we will present a version of the algorithm that takes advantage of the available log-likelihood function in LLMs and constructs the empirical joint and product distributions in a slightly different way. As we will show in the experiments section, $n=2$ is sufficient to have an effective hallucination detection method for the benchmarks that we consider.
The bias introduced by $(\gamma_{1},\gamma_{2})$ in the last equation allows us to rigorously bound the error in estimating $I(\mu)$ via $\widehat{I}_{k}(\gamma_{1},\gamma_{2})$, which is explored next. In particular, in Theorem 4.6 we prove a high-probability lower bound on $I(\mu)$ in terms of $\widehat{I}_{k}$. The core of controlling the estimation error is in accounting for the missing mass, or in other words, how much of $\mu$ we miss out by only observing a finite sample. In Appendix E, we present a more complete discussion and the proof of the bound on the estimation error for mutual information. Here we adapt this result to our particular case.
@@
Using this quantity, we are ready to present a non-asymptotic bound on the estimation error, which depends on the estimator $\widehat{I}_{k}(\gamma_{1},\gamma_{2})$, the expected missing mass, and the sample size:
@@
Suppose that $\widehat{I}_{k}(\gamma_{1},\gamma_{2})$ is given by Algorithm 1, and assume that ${\mathcal{X}}$ is finite. For $\gamma_{1}=1/(k\,|{\mathcal{X}}^{n}|)$, and $\gamma_{2}$ satisfying $\gamma_{2}\geq\gamma_{1}+n(1-Z)$, with probability at least $1-\delta$, we have
$I(\mu)\geq(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma_{1},\gamma_{2})-\left(\frac{1}{k}+(1+n\,\ln\big{(}1+k\,|{\mathcal{X}}|)\big{)}\,\varepsilon_{k}\right)$
@@
Furthermore, given $\delta_{\mathrm{supp}}\in[0,1)$, let $\tilde{\mathcal{X}}\subseteq{\mathcal{X}}^{n}$ such that $\mu(\tilde{\mathcal{X}})\geq 1-\delta_{\mathrm{supp}}$. Then, for $\gamma_{1}=1/(k\,|\tilde{\mathcal{X}}|)$, and $\gamma_{2}$ satisfying $\gamma_{2}\geq\gamma_{1}+n(1-Z)$, with probability at least $1-\delta$, we have
$I(\mu)\geq(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma_{1},\gamma_{2})-\left(\frac{1}{k}+(1+\ln\big{(}1+k\,|\tilde{\mathcal{X}}|)\big{)}\left(\delta_{\mathrm{supp}}+\varepsilon_{k}\right)\right)~{}.$
The theorem is a corollary of Theorem E.4 shown in Appendix E. Note that in Theorem 4.6 we consider two bounds. The first one is pessimistic in the sense that it does not expect that the samples carry much information about the support, and it is most suitable in situations where we expect $\mu$ to be spread out (uniformly) across its entire support. The price of not having samples covering the whole support in this case is a factor $n\ln|{\mathcal{X}}|$ appearing in the bound. For example, in case of a language model with $10,000$ tokens, considering all possible strings of length $T$ tokens yields $n\ln|{\mathcal{X}}|=n\,T\ln(10000)$, and so
$I(\mu)\geq(1-\varepsilon_{k})\,\widehat{I}_{k}(\gamma_{1},\gamma_{2})-\left(\frac{1}{k}+(1+n\,T\ln\big{(}1+k\,\ln(10000))\big{)}\,\varepsilon_{k}\right)~{}.$
Arguably, in practice, such situations are rare, as in natural languages we will not encounter all possible strings. To this end, we consider an optimistic scenario where the effective support of $\mu$, denoted by $\tilde{\mathcal{X}}$, is small with high probability. In this case, we can replace the size of the support for strings of length $n$, $|{\mathcal{X}}|^{n}$, in the first bound with the effective support size $|\tilde{\mathcal{X}}|$, and we only pay essentially a factor $\ln(1+k|\tilde{\mathcal{X}}|)$ instead of $n\ln(1+k|{\mathcal{X}}|)$. In case the effective sample size is only polynomial in $n$, this leads to an exponential reduction in $n$ for the second term in the bounds. In fact, in Section E.4 we demonstrate some empirical evidence that on two question-answering benchmarks, $|\tilde{\mathcal{X}}|$ rarely exceeds $\approx 100$ with $\mu(\tilde{\mathcal{X}})\geq 0.95$, while sampling responses from an LLM given a query.